# Non-existence of Ulrich modules over Cohen-Macaulay local rings

### v1

@@ 4. Examples: Complete intersections
For $k=\mathbb{Q}$, this claim can be verified using, for instance, Macaulay2. Here is a sketch of an argument for general $k$: Since $x$ is a parameter in $S$, it is not a zero-divisor, so it suffices to check that the ring $S[x^{-1}]$ is a domain. One has
@@
It thus suffices to verify that $A$ is a domain and $x^{3}$ does not have a square-root in the fraction field of $A$.
@@
Finally, if $x^{3}$ has a square-root in the fraction field of $A$, then so is $x$. Thus there exist nonzero elements $a,b\in A$ such that $a^{2}x=b^{2}$. Since $A$ can be standard graded, with degrees of $x,y,z$ equal to $1$, this is not possible since the highest degree term in $a^{2}x$ has odd degree while the highest degree term in $b^{2}$ has even degree.
@@
and this is interesting, for it is closely connected to Yhee’s construction [21] of domains (not Cohen-Macaulay) that have no Ulrich modules.
Indeed, according to Macaulay2, the ring $S$ has three minimal prime ideals. Evidently, for any of minimal prime $\mathfrak{p}$ the domain $S/\mathfrak{p}$ (or rather it localization at the maximal ideal) cannot have Ulrich modules, for any Ulrich module over it would be Ulrich over $S_{\mathfrak{m}}$. Again according to Macaulay2, one of the prime ideals is
@@
This subring on the right differ from the subring $k+(u,v)^{2}$ considered in [21], as discussed in the previous section, only in containing $u^{2}v,uv^{2}$, and one can argue as in [21] to deduce that $S/\mathfrak{p}$ does not have Ulrich modules.
@@
$e(S_{\mathfrak{m}})$ $=\operatorname{length}\left(\frac{S}{(x_{11},\dots,x_{nn})}\right)$
$=\operatorname{length}\left(\frac{k[s_{1},\dots,s_{n},x_{ij}\mid 1\leq i<j\leq n]}{(s_{1}^{2},\dots,s_{n}^{2},s_{i}s_{j}-x_{ij}^{3}\mid 1\leq i<j\leq n)}\right)$
@@
where the last equality follows because the ring in question is a finite free extension of $k[s_{1},\dots,s_{n}]/(s_{1}^{2},\dots,s_{n}^{2})$ of degree $3^{\binom{n}{2}}$. Note that $S_{\mathfrak{m}}$ is far from a strict complete intersection, as the product of the $\mathfrak{m}$-adic orders of the defining equations is $2^{\binom{n+1}{2}}$.
One thus gets a family of complete intersections with no Ulrich modules, and for which the difference between the orderand its multiplicity is arbitrarily large.
@@ 5. Examples: Gorenstein normal domains
Let $C$ be a smooth non-hyperelliptic curve of genus $g\geq 3$ (e.g. a smooth plane quartic), and let $\mathcal{L}$ be a non-effective theta characteristic on $C$. In other words,
$\mathcal{L}^{\otimes 2}\cong\omega_{C}\quad\text{and}\quad\Gamma^{0}(C,\mathcal{L})=0\,.$
It is a classical result [16] (see also [1]) that non-effective theta characteristic exists on any smooth curve. For example, one can take
$C=\operatorname{Proj}\frac{\mathbb{C}[x,y,z]}{(x^{4}+y^{4}+z^{4})}\,.$
There is an ample line bundle $\mathcal{L}$ on $C$ such that $\mathcal{L}^{2}=\omega_{C}$ and $\Gamma(C,\mathcal{L})=0$. In fact, for any plane quartic, there are $642^{2g}=$ 64 theta characteristics, but only $28$ of them are effective (corresponding to the $28$ bitangents); see [9, Exercise IV, 2.3(h)].
@@
To begin with $S$ is normal because it is the section ring of smooth projective variety, with respect to an ample invertible sheaf; see [18, Section 1]. It is also Gorenstein by [13, Proposition 3.14(4)]. These conditions are then inherited by $S_{\mathfrak{m}}$ and its completion. For the rest, we show that the conditions in Corollary 3.2 are satisfied with $a=2$.
Indeed, since $C$ is not hyperelliptic, the canonical ring $\operatorname{S}(\omega_{C})$ is generated in degree 1 by a classical theorem of Max Noether. As $\mathcal{L}^{\otimes 2}\cong\omega_{C}$, we deduce that the multiplication map $\mathrm{Sym}^{m}(S_{2})\to S_{2m}$ is surjective for all integer $m$, hence $S_{2}\otimes_{k}S_{j}\to S_{2+j}$ is also surjective for all even integer $j\geq 2$. On the other hand, the line bundle $\mathcal{L}^{\otimes 3}$ on $C$ is Castelnuovo-Mumford 00-regular (with respect to the very ample line bundle $\omega_{C}$). To see this, it suffices to note that $\chi(\mathcal{L})=0$ by Riemann-Roch, and hence
@@
By [14, Theorem 1.8.3], it follows that the multiplication map $S_{3}\otimes\mathrm{Sym}^{m}(S_{2})\to S_{2m+3}$ is also surjective for all integer $m$, which in turn implies the surjectivity of $S_{2}\otimes_{k}S_{j}\to S_{2+j}$ for all odd integer $j\geq 3$. This verifies the condition (2) from Corollary 3.2, while condition (1) is clear from the choice of $\mathcal{L}$. ∎
@@
Remark 5.3.
The ring $S_{\mathfrak{m}}$ in Theorem 5.1 is not a rational singularity. In fact, rational surface singularities have minimal multiplicities (that is to say, the multiplicity is one less than the embedding dimension, see [17, Theorem on page 94]), and such rings admit Ulrich modules, see [2, Proposition 2.5]. It is an interesting question to find an example of a rational singularity with no Ulrich modules.
@@ Acknowledgements.
Many thanks to Kevin Tucker for helpful conversations concerning this work. The authors were partly supported by National Science Foundation (NSF) grants DMS-2001368 (SBI); DMS-2302430 (LM); DMS-1952366 (LM); DMS-2200732 (MW); DMS-2240926 (ZZ), and DMS-2234736 (ZZ). Ma was also partly supported by a fellowship from the Sloan Foundation and by a grant from the Institute for Advanced Study School of Mathematics. Zhuang was also partly supported by a Clay research fellowship, as well as a Sloan fellowship.

### v2

@@ 4. Examples: Complete intersections
For $k=\mathbb{Q}$, this claim can be verified using, for instance, Macaulay2 [7]. Here is a sketch of an argument for general $k$: Since $x$ is a parameter in $S$, it is not a zero-divisor, so it suffices to check that the ring $S[x^{-1}]$ is a domain. One has
@@
It thus suffices to verify that $A$ is a domain and $x^{3}$ does not have a square root in the fraction field of $A$.
@@
Finally, if $x^{3}$ has a square root in the fraction field of $A$, then so is $x$. Thus there exist nonzero elements $a,b\in A$ such that $a^{2}x=b^{2}$. Since $A$ can be standard graded, with degrees of $x,y,z$ equal to $1$, this is not possible since the highest degree term in $a^{2}x$ has odd degree while the highest degree term in $b^{2}$ has even degree.
@@
and this is closely connected to Yhee’s construction [23] of domains (not Cohen-Macaulay) that have no Ulrich modules.
Indeed, according to Macaulay2 [7], the ring $S$ has three minimal prime ideals. Evidently, for any of minimal prime $\mathfrak{p}$ the domain $S/\mathfrak{p}$ (or rather it localization at the maximal ideal) cannot have Ulrich modules, for any Ulrich module over it would be Ulrich over $S_{\mathfrak{m}}$. Again according to Macaulay2 [7], one of the prime ideals is
@@
This subring differs from the subring $k+(u,v)^{2}$ considered in [23], as discussed in the previous section, only in containing $u^{2}v,uv^{2}$, and one can argue as in [23] (or as in the previous section) to deduce that $S/\mathfrak{p}$ does not have Ulrich modules.
@@
$e(S_{\mathfrak{m}})$ $=\operatorname{length}_{S}\left(\frac{S}{(x_{11},\dots,x_{nn})}\right)$
$=\operatorname{length}_{S}\left(\frac{k[s_{1},\dots,s_{n},x_{ij}\mid 1\leq i<j\leq n]}{(s_{1}^{2},\dots,s_{n}^{2},s_{i}s_{j}-x_{ij}^{3}\mid 1\leq i<j\leq n)}\right)$
@@
where the last equality follows because the ring in question is a finite free extension of $k[s_{1},\dots,s_{n}]/(s_{1}^{2},\dots,s_{n}^{2})$ of degree $3^{\binom{n}{2}}$. Note that $S_{\mathfrak{m}}$ is not a strict complete intersection, as the product of the $\mathfrak{m}$-adic orders of the defining equations is $2^{\binom{n+1}{2}}$. One thus gets a family of complete intersections with no Ulrich modules, and for which the difference between the orders and multiplicity is arbitrarily large.
@@ 5. Examples: Gorenstein normal domains
Let $C$ be a smooth non-hyperelliptic curve of genus $g\geq 3$ (for example, a smooth plane quartic) over an algebraically closed field $k$. For simplicity, we assume that char$(k)=0$. Let $\mathcal{L}$ be a non-effective theta characteristic on $C$; in other words
$\mathcal{L}^{\otimes 2}\cong\omega_{C}\quad\text{and}\quad\Gamma(C,\mathcal{L})=0\,.$
It is a classical result [18] (see also [1]) that non-effective theta characteristic exists on any smooth curve. Consider, for example, the curve
$\operatorname{Proj}\frac{\mathbb{C}[x,y,z]}{(x^{4}+y^{4}+z^{4})}\,.$
For any plane quartic, there are $642^{2g}=$ 64 theta characteristics, but only $28$ of them are effective (corresponding to the $28$ bitangents); see [10, Exercise IV, 2.3(h)].
@@
To begin with $S$ is normal because it is the section ring of smooth projective variety, with respect to an ample invertible sheaf; see [20, Section 1]. It is also Gorenstein by [15, Proposition 3.14(4)]. These conditions are then inherited by $S_{\mathfrak{m}}$ and its completion. For the rest, we show that the conditions in Corollary 3.3 are satisfied with $a=2$.
Indeed, since $C$ is not hyperelliptic, the canonical ring $\operatorname{S}(C,\omega_{C})$ is generated in degree 1 by a classical theorem of Max Noether. As $\mathcal{L}^{\otimes 2}\cong\omega_{C}$, we deduce that the multiplication map $\mathrm{Sym}^{m}(S_{2})\to S_{2m}$ is surjective for all integers $m$, hence $S_{2}\otimes_{k}S_{j}\to S_{2+j}$ is also surjective for all even integers $j\geq 2$. On the other hand, the line bundle $\mathcal{L}^{\otimes 3}$ on $C$ is Castelnuovo-Mumford 00-regular (with respect to the very ample line bundle $\omega_{C}$). To see this, it suffices to note that $\chi(\mathcal{L})=0$ by Riemann-Roch, and hence
@@
By [16, Theorem 1.8.3], it follows that the multiplication map $S_{3}\otimes\mathrm{Sym}^{m}(S_{2})\to S_{2m+3}$ is also surjective for all integers $m$, which in turn implies the surjectivity of $S_{2}\otimes_{k}S_{j}\to S_{2+j}$ for all odd integers $j\geq 3$. This verifies the condition (2) from Corollary 3.3, while condition (1) is clear from the choice of $\mathcal{L}$. ∎
@@
The ring $S_{\mathfrak{m}}$ in Theorem 5.1 is also not a rational singularity. In fact, rational surface singularities have minimal multiplicities—see [19, Theorem on page 94]—and such rings admit Ulrich modules, see [2, Proposition 2.5]. It is an interesting question whether there exist rational singularities with no Ulrich modules.
@@ Acknowledgements.
Many thanks to Kevin Tucker for helpful conversations concerning this work, and a referee for their suggestions. The authors were partly supported by National Science Foundation (NSF) grants DMS-2001368 (SBI); DMS-2302430 (LM); DMS-1952366 (LM); DMS-2200732 (MW); DMS-2240926 (ZZ), and DMS-2234736 (ZZ). Ma was also partly supported by a Sloan fellowship and by a grant from the Institute for Advanced Study School of Mathematics. Zhuang was also partly supported by a Clay research fellowship, as well as a Sloan fellowship.